A richly flavored coffee sells for $7 per pound and a hazelnut coffee sells for $4 per pound. The store owner decided to sell a blend of the two coffees. How much of the richly flavored coffee should she mix with 12 pounds of the hazelnut coffee to get a mixture that sells for $6 per pound?
Answer by richwmiller(17219)

Believe it or not, not every problem uses quadratic equations.
7r+4*12=6*(r+12)
r=24
check
7*24+4*12=6*36
168+48=216
ok